A particle with a charge of 34 μC moves with a speed of 75 m/s in the positive x direction. The magnetic field in this region of space has a component of 0.42 T in the positive y direction, and a component of 0.87 T in the positive z direction.What is the magnitude of the magnetic force on the particle?What is the direction of the magnetic force on the particle? (Find the angle measured from the positive z-axis toward the negative y-axis in the yz-plane.)
Answer #1

here,

the charge on particle , q = 34 uC = 34 * 10^-6 C

the speed of particle , v = 75 m/s i

the magnetic feild ,B = 0.42 j T + 0.87 k T

the magnetic force , F = q * ( v X B)

F = 34 * 10-6 * ( 75 i X ( 0.42 j + 0.87 k))

F = (1.07 k - 2.22 j) * 10^-3 N

the magnitude of the magnetic force on the particle , |F| = sqrt(1.07^2 + 2.22^2) * 10^-3 N

|F| = 2.46 * 10^-3 N

the direction of magnetic force , theta = arctan(2.22 /1.07) = 64.3 degree